# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v2

@@ IV-D Ablation Studies
To thoroughly evaluate the effectiveness of our proposed method, we conduct an ablation study on the University-1652 dataset. We systematically assess the impact of each individual module and investigate their contributions under different configurations. The results are presented in Table LABEL:Ablation_Studies.
ConvNeXt-Tiny Backbone. The Backbone refers to the untrained ConvNeXt-Tiny network used as a direct feature extractor. As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, achieving an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone Satellite and Satellite Drone, respectively. These results indicate that the backbone alone lacks the ability to extract discriminative cross-view features.
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the dual-path contrastive learning baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between UAV and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with the contrastive learning baseline, incorporating DHML leads to significant performance improvements under both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in UVGL scenarios and enhances intra-view feature representation and discrimination.
@@
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to within-view information. To address this limitation, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%. These results further validate the effectiveness of ICEL in enhancing cross-view consistency for UVGL, thereby contributing to overall performance improvement.
Pseudo-Label Enhancement (PLE): Refining Feature Learning with Robust Labels. To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability. After incorporating PLE, our final model achieves state-of-the-art performance, with R@1 increasing to 90.17% and 95.01%, and AP reaching 91.67% and 88.95%. This confirms that refining pseudo-label quality significantly enhances the final model’s effectiveness in cross-geographic view matching.
The ablation study systematically evaluates the contribution of each component in the proposed self-supervised framework. Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline. On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement. Experimental results show that our method significantly enhances the learning of discriminative and view-consistent representations under self-supervised learning.
@@ IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the our method, We perform feature space (t-SNE[55] map) and similarity distribution visualization for our method, as presented in Fig 4.Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct, indicating better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our method in UVGL.
@@
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts across epochs for three cross-view datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater UAV image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (1,200 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates the ability of our method to progressively learn discriminative representations.
@@
Retrieval results. To demonstrate the effectiveness of our method, we present retrieval results on the University-1652 and DenseUAV dataset in Fig. 6 and Fig. 7. Green borders indicate correct matches, while red borders denote errors, showing that our method significantly improves retrieval accuracy.
In the University-1652 dataset, the dronesatellite retrieval task presents drone-view images as queries and retrieves corresponding satellite-view images. Since each location is associated with a single satellite image, our method consistently achieves Top-1 accuracy, demonstrating its high effectiveness in cross-view matching. In the satellitedrone setting, our method accurately retrieves the corresponding drone-view images given satellite queries, further showcasing its strong cross-view localization capability and reliable feature alignment across perspectives.
@@
In the DenseUAV dataset, the retrieval task becomes significantly more challenging due to dense spatial sampling, multi-scale satellite imagery, and temporal dynamics, all of which exacerbate the difficulty of cross-view representation learning. In the Drone Satellite retrieval setting, despite variations in UAV altitudes and satellite image resolutions, our method effectively captures spatial-geographic cues, achieving high retrieval accuracy while substantially reducing mismatches. Compared to University-1652, DenseUAV provides a more realistic and complex benchmark for UAV-based geo-localization. Under this setting, our method exhibits strong cross-view retrieval performance, demonstrating robust feature discrimination in densely distributed urban regions.
@@ V CONCLUSION
This paper investigates a highly valuable and challenging task namely self-supervised UAV-view cross-view geo-localization, aiming to reduce the reliance on paired UAV–satellite images. To this end, we propose a self-supervised UVGL method, which employs a shallow backbone with only 9 layers for feature extraction. Pseudo labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module. Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across views. To improve the quality and robustness of pseudo labels, we introduce a bidirectional feature consistency-based Pseudo Label Enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, even surpassing several fully supervised methods in certain scenarios, without relying on paired UAV–satellite images. These results validate the effectiveness and generalization capability of our method, highlighting its potential for real-world UVGL applications.

### v3

@@ IV-D Ablation Studies
To thoroughly evaluate the effectiveness of our proposed method, we perform an ablation study on the University-1652 dataset. We systematically assess the impact of each individual module and investigate their contributions under different configurations. The results are presented in Table LABEL:Ablation_Studies.
ConvNeXt-Tiny Backbone. The backbone refers to the untrained ConvNeXt-Tiny network used as a direct feature extractor. As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, and obtains an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone Satellite and Satellite Drone, respectively. These results indicate that the backbone alone lacks the ability to extract discriminative cross-view features.
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between drone and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with the baseline, the incorporation of DHML leads to significant performance improvements in both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in DVGL scenarios and enhances intra-view feature representation and discrimination.
@@
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to intra-view information. To address this limitation, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%. These results validate the effectiveness of ICEL in enhancing cross-view consistency for DVGL.
Pseudo-Label Enhancement Refining Feature Learning with Robust Labels. To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability. After incorporating PLE, our final model achieves state-of-the-art performance, with R@1 increasing to 90.17% and 95.01%, and AP reaching 91.67% and 88.95%. This confirms that refining pseudo-label quality significantly enhances the final model’s effectiveness in cross-geographic view matching.
The ablation study systematically evaluates the contribution of each component in our method. Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline. On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement. Experimental results show that our method enhances the learning of discriminative and view-consistent representations under self-supervised learning.
@@ IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the method, we perform a feature space and similarity distribution visualization for our method, as presented in Fig. 4. Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct that indicates better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our method.
@@
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts in each epoch for three datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater drone image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (120 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates the ability of our method to learn discriminative representations progressively.
@@
Retrieval results. To demonstrate the effectiveness of our method, we present retrieval results on the University-1652 and DenseUAV datasets in Fig. 6 and Fig. 7. Green borders indicate correct matches, while red borders denote errors, showing that our method significantly improves retrieval accuracy. In the University-1652 dataset, the dronesatellite retrieval task uses drone-view images as queries to retrieve corresponding satellite images. Our method achieves consistent Top-1 accuracy to highlight its effectiveness in cross-view matching. In the satellitedrone setting, our method also accurately retrieves drone-view images from satellite queries that demonstrates strong cross-view retrieval ability.
@@
Compared with University-1652, the DenseUAV dataset introduces more realistic and complex challenges, such as dense spatial sampling, multi-scale resolutions, and temporal dynamics. In the drone satellite retrieval setting, despite the high risk of confusion caused by dense spatial sampling, our method effectively captures spatial-geographic cues and achieves high retrieval accuracy, demonstrating robust cross-view performance under realistic and complex conditions.
@@ V CONCLUSION
This paper investigates a highly valuable and challenging DVGL task to reduce reliance on paired drone–satellite images. To this end, we propose a self-supervised learning method that employs a shallow backbone with only 9 layers for feature extraction. Pseudo-labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module. Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across-views. To improve the quality and robustness of pseudo-labels, we introduce a pseudo-label enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, and even surpasses several fully supervised methods in certain scenarios without relying on paired drone–satellite images. These results validate the effectiveness and generalization capability of our method and verify the potential for real-world DVGL applications.